# Fair Generation without Unfair Distortions: Debiasing Text-to-Image Generation with Entanglement-Free Attention

### v1

@@ 5.1 Effect of regularization and segmentation mask
Table 3: Effectiveness of the regularization loss and the segmentation mask in bias mitigation and non-target attribute preservation. P. and seg. are abbreviations of the preservation and segmentation, respectively.
@@
When training without segmentation masks, we exclude $\mathcal{L}_{\text{reg}}^{\text{cf}}$ in training and apply $\mathcal{L}_{\text{recon}}$ to the entire image. Since there is no constraint limiting target attribute enhancement, we observe a slight improvement in bias mitigation performance. However, without segmentation masks to distinguish human and non-human regions during training, EFA struggles to localize them precisely, leading to reduced preservation of non-target attributes at inference. Nonetheless, EFA still outperforms baselines even without segmentation masks. We conjecture that this is because counterfactual attributes are primarily activated in bias-relevant regions due to their semantic similarity in cross-attention, which encourages EFA to enhance the target attribute in those regions while ignoring the counterfactual prompt’s influence. This allows EFA to approximate the localization of bias-relevant regions without relying on fine-grained masks of the target attribute. Also, when $\mathcal{L}_{\text{reg}}$ is not applied, the non-target attribute preservation performance deteriorates, even though the bias mitigation performance remains similar. These results highlight the necessity of the regularization mechanisms to preserve non-target attributes across diverse scenarios where the target attribute appears at different extents in the input.
@@ 5.2 Analyzing the influence of resolution in EFA
Table 4: The impact of input resolution employed by EFA mainly affects non-target preserving while having little effect on bias. P. denotes preservation.
@@ 5.3 Attention map visualization
In Fig. 5, we visualize the attention maps predicted by EFA when the target attribute is enhanced. EFA assigns high values to the human region, effectively enabling gender transformation while minimally affecting accessories, clothing style, and background details, thereby preserving the original model’s generative capacity.
@@ 6 Conclusion
In this work, we tackled attribute entanglement in diffusion-based T2I models, where bias mitigation often distorts non-target attributes, causing unintended distribution shifts. Our proposed method, EFA, effectively mitigates human-centric biases while preserving non-target attributes. During training, EFA is guided to attend to regions with high semantic relevance to counterfactual prompts, constrained by human segmentation masks to focus on human regions. At inference time, EFA randomly samples a target attribute and alters the cross-attention in selected layers to ensure a fair distribution of target attributes in the outputs. Extensive experiments show that EFA outperforms prior methods in mitigating bias while preserving the original model’s output distribution and generative capacity.
@@ 7 Implementation details
Architecture. The spatial information is crucial for distinguishing between target and non-target attributes. Therefore, we adopt convolutional networks as our backbone, given their ability to capture and process spatial structural information. Specifically, AP consists of three 2D convolutional layers and two SiLU activation functions. AP predicts attention values for token embeddings of target attributes for each attention head, and the token embeddings of the target attribute include special tokens (i.e., start-of-sequence and end-of-sequence tokens). The target attribute and user prompt are separately processed during the text embedding stage.
Training and Inference. For experiments, NVIDIA RTX 3090, A6000, or V100 GPUs are employed. The models for gender, race, and their intersectional biases were trained for 20 epochs, while the model incorporating age as an attribute was optimized for 10 epochs. Hyperparameters were selected based on validation set performance. When training multiple target attributes simultaneously, we adjusted individual hyperparameters as needed to ensure optimal learning. During inference, we set the timestep for image generation to 50 for all baselines.

### v2

@@ 5.1 Effect of regularization and segmentation mask
Table 3: Effectiveness of the regularization loss and the segmentation mask in bias mitigation and non-target attribute preservation. P. and seg. indicate preservation and segmentation, respectively.
@@
When training without segmentation masks, we exclude $\mathcal{L}_{\text{reg}}^{\text{cf}}$ in training and apply $\mathcal{L}_{\text{recon}}$ to the entire image. Since there is no constraint limiting target attribute enhancement, we observe a slight improvement in bias mitigation performance. However, without segmentation masks to distinguish human and non-human regions during training, EFA struggles to localize them precisely, leading to reduced preservation of non-target attributes at inference. Nonetheless, EFA still outperforms baselines even without segmentation masks. We conjecture that this is because counterfactual attributes are primarily activated in bias-relevant regions due to their semantic similarity in cross-attention, which encourages EFA to enhance the target attribute in those regions while ignoring the counterfactual prompt’s influence. This allows EFA to approximate the localization of bias-relevant regions without relying on fine-grained masks of the target attribute. Also, when $\mathcal{L}_{\text{reg}}^{\text{trg}}$ is not applied, the non-target attribute preservation performance deteriorates, even though the bias mitigation performance remains similar. These results highlight the necessity of the regularization mechanisms to preserve non-target attributes across diverse scenarios where the target attribute appears at different extents in the input.
@@ 5.2 Analyzing the influence of resolution in EFA
Table 4: Impact of input resolution employed by EFA. Extending EFA to higher-resolution layers compromises the preservation of non-target attributes. P. denotes preservation.
@@ 5.3 Attention map visualization
In Fig. 5, we visualize the attention maps predicted by EFA when the target attribute is enhanced. EFA assigns high values to the human region, effectively enabling gender transformation while preserving the generation capability of the original model, as the impact on accessories, clothing style, and background details remains minimal.
@@ 6 Conclusion
In this work, we addressed attribute entanglement in diffusion-based T2I models, where bias mitigation often distorts non-target attributes, leading to unintended distribution shifts. We propose EFA, which mitigates human-centric biases while preserving non-target attributes. During training, EFA is guided to attend regions with high semantic relevance to counterfactual prompts, constrained by human segmentation masks. At inference, EFA randomly samples a target attribute and alters the cross-attention in selected layers to ensure a fair distribution of target attributes in the outputs. Experimental results verify that EFA outperforms prior methods in mitigating bias while preserving the original model’s output distribution and generative capacity.
Acknowledgments
This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2021-II212068, Artificial Intelligence Innovation Hub), and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(No. 2022R1A5A7083908).
@@ 7 Implementation details
Architecture. The spatial information is crucial for distinguishing between target and non-target attributes. Therefore, we adopt convolutional networks as our backbone, given their ability to capture and process spatial structural information. Specifically, AP consists of three 2D convolutional layers and two SiLU activation functions. AP predicts attention values for token embeddings of target attributes for each attention head, and the token embeddings of the target attribute include special tokens (i.e., start-of-sequence and end-of-sequence tokens). The target attribute and user prompt are separately processed during the text embedding stage. To leverage rich semantic information, EFA is applied to selected up-block layers where the input resolution is low. Specifically, the input to EFA is given by $\textbf{z}_{t}\in\mathbb{R}^{B\times S\times D}$, where $B$, $S$, and $D$ denote the batch size, spatial resolution ($S=16\times 16$), and feature dimension, respectively. Note that applying EFA does not change the input or output shapes of the cross-attention module, since the target attribute features are added only to the key and value matrices while leaving the overall architecture intact.
Training and Inference. For experiments, NVIDIA RTX 3090, A6000, or V100 GPUs are employed. The models for gender, race, and their intersectional biases were trained for 20 epochs, while the model incorporating age as an attribute was optimized for 10 epochs. Hyperparameters were selected based on validation set performance. When training with multiple target attributes, we independently tuned the hyperparameters for each attribute to ensure optimal learning. During inference, we set the timestep for image generation to 50 for all baselines.
@@ 10 Simultaneous mitigation of gender, racial, and age biases
11 Efficiency Analysis
We evaluated EFA that simultaneously mitigates gender, race, and age biases on an NVIDIA A100 GPU. The number of parameters increased by 9.4% compared to the original SD (1066M $\to$ 1166M), and the additional inference time is minimal at 0.29 seconds (3.69s $\to$ 3.98s). This overhead could potentially be reduced through implementation optimizations and exploration of lightweight variants of EFA for deployment in resource-constrained environments.
@@ 14 Additional quantitative and qualitative results
Figure 7: Images generated by SD and our method in complex scenes. (a) and (b) show the results of SD and our method that mitigates racial bias, respectively. The last row indicates the target attributes, and the corresponding aligned images were generated using the same random seed. Our method (EFA) successfully preserves non-target attributes while generating complex scenes.
@@
Complex Scenes. Fig. 7 presents results on complex scenes containing multiple individuals, diverse objects, and cultural contexts. Our EFA model for mitigating race bias successfully introduces diverse racial attributes while preserving non-target attributes and maintaining complex scene structure. However, in cases where the SD fails to generate clear facial features (e.g., the last column of Fig. 7), our method struggles to accurately localize and interpret the facial region, thereby limiting its ability to address race-related cues.
15 Limitations and Future Work
Assuming a fixed type of bias may limit the general applicability of debiasing approaches in real-world scenarios. Recent advances in automatic bias identification for generative models [6] offer orthogonal solutions that could complement our work. Integrating such techniques into our framework presents a promising direction for future research. Furthermore, while our work primarily focuses on human-centric biases, a more systematic investigation into non-human-centric biases remains a valuable topic for further exploration.